TYPE A PROBLEM.

CHOOSE ONE INTEGER FROM THE FOLLOWING SET: -5 ≤ X ≤ 5. STATE THE INTEGER YOU CHOSE.

WHEN USING LONG DIVISION, WHAT BINOMIAL WOULD YOU DIVIDE BY TO DETERMINE IF THE INTEGER YOU CHOSE IS A ZERO FOR YOUR POLYNOMIAL?

DO THE LONG DIVISION

WRITE THE ANSWER TO THE LONG DIVISION PROBLEM IN THE FORM QUOTIENT + REMAINDER/DIVISOR


1
Expert's answer
2021-06-15T18:13:26-0400

We have to perform long division

Let's take an equition

"9x^2+8x^2-4x^2-x+7"

We have to choose integer from the following set ":-5\leq X\leq 5"

"X=4"

Then, "x-4=0"

Now

"19x^3+84x^332+1327"

"\sqrt[x-4]{9x^2+8x^2-4x^2-x+7}"

"(19x^4-76x^3)+\\84x^3-4x^2-x+7\\84x^3-336x^2" +

"332x^2-x+7\\332x^2-1328x" +

"1327x+7\\1327x-5308=5315"

Hence

"Quotient+\frac{remainder}{divisor}"

So, "19x^3+84x^2+332x+1327+\frac{5315}{x-4}"

If I choose 0 integer for long Division then x=0

X-0=0 is the binomial

but I chose x=4 for my equition.